OpenStudy (anonymous):

One zero for x^2−10x+169=0 is x=5+12i. Find the second zero for x^2−10x+169. Explain how you got your answer.

OpenStudy (anonymous):

i will fan and Medal

OpenStudy (anonymous):

a*x^2+b*x+c=0 D= b^2-4ac D= 100 - 169*4 D<0